HIGHEST RESOLUTION FOURIER TRANSFORM INFRARED SPECTROSCOPY WITH AN ELEVEN CHAMBER BRUKER INTERFEROMETER AT THE SWISS SYNCHROTRON

\underline{S. ALBERT}, K.K. ALBERT, M. QUACK, PHYSICAL CHEMISTRY, ETH ZÜRICH, CH-8093 ZÜRICH, SWITZERLAND; PH. LERCH, L. QUARONI, SWISS LIGHT SOURCE, PAUL-SCHERRER-INSTITUTE, CH-5232 VILLIGEN, SWITZERLAND ; A. KEENS, BRUKER OP-TICS, D-76275 ETTLINGEN, GERMANY.

We have interfaced a newly constructed eleven chamber interferometer, the ETH-SLS Bruker IFS 125 HR prototype 2009, to the infrared port available at the Swiss Light Source (SLS), located at the Paul-Scherrer-Institute. The Maximum Optical Path Difference (MOPD) of this spectrometer is 11.70 m allowing for a best theoretical unapodized resolution of 0.00053 cm^{−1} (18 MHz). The ETH-SLS Bruker spectrometer is a further development of our nine chamber interferometer Bruker IFS 120/125 Zurich prototype 2001^{a}^{,}^{b} which has an MOPD of 9.4 m and unapodized resolution of 0.00068 cm^{−1} (23 MHz). We present spectra of CO and pyrimidine (C_{4}H_{4}N_{2}) as examples to illustrate the improved resolution. Due to the high brightness of the synchrotron source the signal-to-noise ratio is effectively 5 to 20 times better than that of conventional thermal sources in the spectral region between 180 and 900 cm^{−1} (6-28 THz). We present examples of pyrimidine (C_{4}H_{4}N_{2}) and CDBrClF spectra in the region 600 to 900 cm^{−1} and of phenol (C_{6}H_{5}OH) and aniline (C_{6}H_{5}NH_{2}) spectra in the region 180 to 350 cm^{−1}. Due to the excellent resolution and the bright synchrotron source we were able to detect the spin statistical weights in the pyrimidine^{c} spectra and found new combination bands in the CDBrClF spectra. We were able to rotationally resolve the torsional c-type band of phenol with ν_{0a} = 309.1141 cm^{−1} and ν_{0b} = 309.5517 cm^{−1} and detect a torsional splitting of 0.4376 cm^{−1} in the v = 1 torsional level. In addition, we were able to rotationally resolve and assign the very weak two torsional b-type subbands of aniline with ν_{0a} = 234.8 cm^{−1} and ν_{0b} = 304.3 cm^{−1} showing the mode selective inversion splitting^{d}.

aS. Albert, K.K. Albert and M. Quack, Trends in Optics and Photonics, 2003, 84, 177.

bS. Albert and M. Quack, ChemPhysChem, 2007, 8, 1271.

cS. Albert and M. Quack, J. Mol. Spectrosc., 2007, 243, 280.

dB. Fehrensen, M. Hippler and M. Quack, Chem. Phys. Lett., 1998, 298, 320.